### v1

@@ 5.4.1 Feature Extraction Module
For Light-VQA+, a significant advancement lies in the method used for extracting the BNF and BCF. Notably, large language models demonstrate superior performance compared to traditional methods in assessing video quality. Model 1 and Model 2 are specifically designed to evaluate the efficacy of CLIP in capturing features. To ensure that the test results remain unaffected by the feature-fusion and regression methods, we also disabled the cross-attention module and weight parameters in both Model 1 and Model 2. The results clearly indicate that CLIP significantly outperform traditional methods in extracting such features from frames. Furthermore, to assess the effectiveness of integrating both local and global perspectives, we developed Model 3 and Model 4. The findings confirm that this combination provides substantial benefits in feature extraction.
5.4.2 Feature Fusion & Regression Module
@@
Extracting features is as crucial as utilizing them effectively. To this end, verifying the efficacy of the cross-attention module is essential. While Model 1,2,3,4 omit the cross-attention fusion modules to isolate and assess the feature extraction capabilities, applying the same strategy in reverse is impractical. This is because our cross-attention modules are specifically designed to integrate the features captured by LLM, not those that are handcrafted. Consequently, Model 5 is crafted to test the effectiveness of the cross-attention module by excluding it from the final model. Additionally, the importance of weights for aligning HVS also warrants verification. Thus, Model 6 excludes these weights during the training process. By comparing the outcomes from Model 5 and Model 6 with the complete setup (i.e., Model 7), it becomes apparent that both the cross-attention and weights for aligning the HVS play significant roles in enhancing the predictive performance of Light-VQA+.
5.5 Refinement for Training Exposure Correction Algorithms
@@
Table 6: The detailed qualitative comparisons of experiments. The videos we test are the original low-light and over-exposed videos from VEC-QA corrected by the trained model. The MOS presented are the average MOS of these corrected videos given by Light-VQA+. The pictures below are the example frames from the corrected videos. [Key: Best; Second Best]
Video Original Videos W/O Light-VQA+ W/ Light-VQA+ W/ Light-VQA+
MOS[OE/LL] 68.1/46.2 69.6/59.8 72.5/63.8 74.2/64.0
Training Dataset None SICE [18]&MSEC [1] SICE [18] MSEC [1]
Over Exposed [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Low Light [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
To demonstrate that our Light-VQA+ can serve as an effective metric for facilitating the development of exposure correction algorithms, we utilized Light-VQA+ to fine-tune a recent exposure correction algorithm, FEC-Net [18]. Specifically, we employed the MSEC [1] and SICE [18] datasets for our experiments. The MSEC dataset comprises a diverse collection of frames, including both low-light and over-exposed samples, whereas the SICE dataset consists of two subsets specifically tailored for low-light enhancement and over-exposed recovery, respectively.
After initial training, we tested the modified FEC-Net on original low-light and over-exposed videos from the VEC-QA dataset and assessed the processed videos using our trained Light-VQA+ model. The training of FEC-Net involves firstly training on the MSEC and SICE datasets, followed by fine-tuning using our Light-VQA+ model. During the fine-tuning phase, the assessments from Light-VQA+ are incorporated as part of the loss function. Experimental results demonstrate that using Light-VQA+ in the loss function enhances the perceptual quality of videos corrected by the algorithm. The experimental results and detailed qualitative comparisons are presented in Table 6. It is evident from these findings that FEC-Net, when trained with Light-VQA+, delivers superior perceptual quality across low-light, over-exposed, and general datasets.
@@ 6 Limitation
Although Light-VQA+ has demonstrated impressive performance on our VEC-QA dataset, there remains potential for further enhancements. Powered by CLIP [40], Light-VQA+ has the capability to not only assess the noise and brightness attributes of video frames, but also to explain the logic behind its predictions. Unfortunately, limitations in available corpora and computational resources have prevented us from realizing this capability fully. In future work, we plan to enable Light-VQA+ to provide explanations for its predictions, which will enhance its utility and transparency. Additionally, as new over-exposure recovery algorithms designed specifically for video become available, we intend to expand our VEC-QA dataset by applying these algorithms to our existing data. This expansion will allow us to continuously update and improve the performance of our Light-VQA+ model.
@@ 7 Conclusion
In this paper, we address the challenge of evaluating the quality of VEC algorithms. We introduce the OEVR-QA dataset, which comprises 2,458 videos, including 205 original over-exposed videos and their corrected versions. Together with the LLVE-QA dataset [8], we have formed the VEC-QA dataset consisting of 4,518 videos. Additionally, we propose Light-VQA+, an advanced VQA model tailored for assessing videos with improper exposure. Specifically, Light-VQA+ leverages the capabilities of CLIP [40] through a set of specially designed vision-language prompts to extract relevant features. These features are integrated with deep-learning-based spatial and temporal information, combining local and global insights for a comprehensive analysis. The fusion of spatial and temporal data is achieved using a cross-attention method [13]. After quality regression, a set of weights are further applied to make our model align with the HSV [4]. Our experimental results confirm the effectiveness of Light-VQA+. We anticipate that Light-VQA+ will serve as a foundational model in the field of VEC algorithm assessment.

### v2

@@ 5.4.1 Feature Extraction Module
For Light-VQA+, a significant advancement lies in the method that extracts the BNF and BCF. Notably, multimodal large language models demonstrate the superior performance compared to traditional methods in assessing video quality. Model 1 and Model 2 are specifically designed to evaluate the efficacy of CLIP in capturing features. To ensure that the test results remain unaffected by the feature-fusion and regression modules, we also disabled the cross-attention module and weight parameters in both Model 1 and Model 2. The results clearly indicate that CLIP significantly outperforms the traditional methods in extracting such features from frames. Furthermore, the results of Model 3 and Model 4 validate the effectiveness of integrating both local and global perspectives.
5.4.2 Feature Fusion & Regression Modules
Extracting features is as crucial as utilizing them effectively. To this end, verifying the efficacy of the cross-attention module is essential.
@@ 5.4.2 Feature Fusion & Regression Modules
To this end, Model 5 is crafted to test the effectiveness of the cross-attention module by excluding it from the final model. Additionally, the importance of HVS weights should also be verified. Thus, Model 6 excludes these HVS weights during the training process. By comparing the outcomes from Model 5 and Model 6 with our full model (i.e., Light-VQA+), it becomes apparent that both the cross-attention and HVS weights play significant roles in enhancing the predictive performance of Light-VQA+.
5.5 Benefits to Exposure Correction Algorithms
@@
Table 6: The detailed qualitative comparisons of experiments. The videos we test are the original low-light and over-exposed videos from VEC-QA corrected by the pre-trained and fine-tuned FEC-Net models. The scores presented are the average of the corrected 100 videos assessed by Light-VQA+. The figures below are the example frames from the corrected videos. [Key: OE: Over Exposed; LL: Low Light; Score: OE/LL; Best; Second Best]
Video Original Pre-trained Fine-tuned
Score 68.1/46.2 69.6/59.8 74.2/64.0
OE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
LL [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
To demonstrate that our Light-VQA+ can serve as an effective metric for facilitating the development of exposure correction algorithms, we utilize the Light-VQA+ to fine-tune a recent exposure correction algorithm, FEC-Net [82]. We leverage the released FEC-Net model that is pre-trained on SICE dataset, which consists of two subsets that collect the low-light and over-exposed images respectively.
Subsequently, the Light-VQA+ is utilized to fine-tune the FEC-Net. During the fine-tuning phase, the result of Light-VQA+ is incorporated as part of the loss function. After training and fine-tuning, we randomly choose 50 over-exposed and 50 low-light original videos from VEC-QA, and use the pre-trained and fine-tuned FEC-Net models to correct them. Finally, we assess the quality of corrected videos.
As shown in Tab. 6, the pre-trained FEC-Net model is capable of correcting the improper exposed videos. With the help of Light-VQA+, the performance of fine-tuned FEC-Net is further improved in exposure correction, which justifies the effectiveness of our quality assessment model.
@@ 6 Limitation
Although Light-VQA+ has demonstrated impressive performance on our VEC-QA dataset, there remains potential for further improvement. It is important to note that the CLIP is not specifically designed for VQA tasks. In general, modifying CLIP through fine-tuning could enhance its effectiveness for this specific application. However, the lack of suitable datasets for the quality assessment task hinders us from fine-tuning the CLIP. In the future, we plan to collect a specific dataset that can support the CLIP fine-tuning.
@@ 7 Conclusion
In this paper, we focus on the issue of evaluating the quality of VEC algorithms. To facilitate our work, we construct a VEC-QA dataset containing 4,518 videos that feature diverse content with various brightness levels. Further, we propose an effective VQA model named Light-VQA+. Concretely, we integrates the exposure-sensitive CLIP-captured features into both spatial and temporal features. After the feature fusion, we perform a quality regression to obtain assessment scores for each video clip. Finally, we apply a weighted average on these results with the HVS weights to obtain the final quality assessment score. Our experimental results confirm the effectiveness of Light-VQA+ as a video quality assessment model for exposure correction. Besides, the Light-VQA+ is capable of boosting the performance of video exposure correction methods. Therefore, we envision that the Light-VQA+ would be widely considered while developing them.
Data Availability. The datasets used for experiments during the current study are all publicly available. The KoNViD-1k is available at https://database.mmsp-kn.de/konvid-1k-database.html. The VDPVE dataset is available at https://github.com/YixuanGao98/VDPVE-VQA-Dataset-for-Perceptual-Video-Enhancement. The YouTube-UGC dataset is available at https://media.withyoutube.com/. The LIVE-VQC is available at https://live.ece.utexas.edu/research/LIVEVQC/index.html. The SICE dataset is available at https://github.com/KevinJ-Huang/FECNet.
Acknowledgement. The work was supported in part by the National Natural Science Foundation of China under Grant 62301310, and in part by the Shanghai Pujiang Program under Grant 22PJ1406800.